As a condition of employment, Fashion Industries applicants must pass a drug test. Of the last 220 applicants, 18 failed the test. (Use z Distribution Table.) a. Develop a 90% confidence interval for the proportion of applicants that fail the test. (Round your answers to 3 decimal places.) For the applicants the confidence interval is between and . b. Would it be reasonable to conclude that more than 12% of the applicants fail the test?
Answer #1

Solution:

a)

Given,

n = 220 ....... Sample size

x = 18 .......no. of successes in the sample

Let \hat p denotes the sample proportion.

  \therefore  \hat p = x/n = 18/220 = 0.0818

Our aim is to construct 90% confidence interval.

\therefore c = 0.90

\therefore\alpha = 1- c = 1- 0.90 = 0.10

\therefore  \alpha/2 = 0.10 \slash 2 = 0.05 and 1- \alpha /2 = 0.950

\therefore2012 = 1.645 (use z table)

Now , the margin of error is given by

E = 2012 *  Vp(1-P)/n

= 1.645 * \sqrt{} [0.0818 *(1 - 0.0818)/220]

= 0.0304

Now the confidence interval is given by

(\hat p - E)  < p < (\hat p + E)

(0.0818   - 0.0304) < p <   (0.0818   + 0.0304)

0.051  < p < 0.112

For the applicants the confidence interval is between 0.051 and  0.112

b)

Claim : More than 12% of the applicants fail the test.

i.e. p > 0.12

Hypothetical value 0.12 is not in the interval. But lower limit 0.051 is not greater than 0.12

So , fail to reject H0 : p \leq 0.12

Answer is

No , it would not be reasonable to conclude that more than 12% of the applicants fail the test.